# FG-CLIP: Fine-Grained Visual and Textual Alignment

### v2

@@ Abstract
Contrastive Language-Image Pre-training (CLIP) excels in multimodal tasks such as image-text retrieval and zero-shot classification but struggles with fine-grained understanding due to its focus on coarse-grained short captions. To address this, we propose Fine-Grained CLIP (FG-CLIP), which enhances fine-grained understanding through three key innovations. First, we leverage large multimodal models to generate 1.6 billion long caption-image pairs for capturing global-level semantic details. Second, a high-quality dataset is constructed with 12 million images and 40 million region-specific bounding boxes aligned with detailed captions to ensure precise, context-rich representations. Third, 10 million hard fine-grained negative samples are incorporated to improve the model’s ability to distinguish subtle semantic differences. We construct a comprehensive dataset, termed FgGRN, by integrating high-quality region-specific annotations with challenging fine-grained negative samples. Corresponding training methods are meticulously designed for these data. Extensive experiments demonstrate that FG-CLIP outperforms the original CLIP and other state-of-the-art methods across various downstream tasks, including fine-grained understanding, open-vocabulary object detection, image-text retrieval, and general multimodal benchmarks. These results highlight FG-CLIP’s effectiveness in capturing fine-grained image details and improving overall model performance. The data, code, and models are available at https://github.com/360CVGroup/FG-CLIP.
@@
Dataset: https://huggingface.co/datasets/qihoo360/FgGRN-12M
@@ 1 Introduction
Third, to further enhance model robustness and discrimination abilities, we introduce a large-scale corpus of 10 million hard fine-grained negative samples. By incorporating these challenging negative samples into the training process, FG-CLIP learns to distinguish subtle differences in semantically similar but distinct pairs, thereby significantly improving its performance across various downstream tasks. We integrate the high-quality visual grounding data and hard fine-grained negative samples as a whole dataset called FgGRN.
@@ 3.2 Curated Dataset
Fine-grained visual Grounding+Recaption+Negative dataset(FgGRN).
For the second stage, we develop a high-quality visual grounding dataset named FgGRN, featuring precise region-specific captions and challenging negative samples. We curate the overall dataset based on GRIT (Peng et al., 2024) images. The process begins with generating detailed image captions using CogVLM2-19B (Hong et al., 2024), ensuring comprehensive and nuanced descriptions that capture the full context of each image. Following (Peng et al., 2024), we then use SpaCy (Honnibal et al., 2020) to parse the captions and extract the referring expressions. Subsequently, the images and referring expressions are fed into the pretrained object detection model, i.e., Yolo-World (Cheng et al., 2024) to obtain the associated bounding boxes. Non-maximum suppression is applied to eliminate overlapping bounding boxes, retaining only those with predicted confidence scores higher than 0.4. This process results in 12 million images and 40 million bounding boxes with fine-grained region captions. We provide examples of the images and their corresponding captions in Appendix A.
@@ 4.1 Implementation Details
In the first stage, we employ 160$\times$×ASCEND 910B NPUs and train on a dataset of 1.6 billion images, each paired with short and long texts. The model is initialized with weights from the original CLIP (Radford et al., 2021). For both ViT-B and ViT-L (Dosovitskiy, 2021) configurations, the batch size per NPU is set to 384. The learnable temperature parameter $\tau$ is initialized to 0.07. We utilize the AdamW optimizer with a learning rate of 1e-4, weight decay of 0.05, $\beta_{1}$ of 0.9, $\beta_{2}$ of 0.98, and warmup steps for the first 200 iterations. The entire training process employs DeepSpeed’s Zero-2 optimization technique and Bfloat16 precision to accelerate training, and the model is trained for one epoch.
@@
In the second stage, we employ 8$\times$×NVIDIA H800 GPUs and train on a dataset of 12 million images. Apart from long and short captions, this dataset includes high-quality visual grounding annotations and hard fine-grained negative samples. The model is initialized with weights obtained from the first stage. The batch size per GPU is set to 512. We employ the AdamW optimizer with a learning rate of 1e-6, weight decay of 0.001, $\beta_{1}$ of 0.9, $\beta_{2}$ of 0.98, and warmup steps for the first 50 iterations. Training acceleration techniques include DeepSpeed’s Zero-2 optimization, CUDA’s TF32 technology, and Bfloat16 precision, and the model is trained for one epoch.
@@ 5 Conclusion
In this work, we introduce Fine-Grained CLIP (FG-CLIP), a novel approach that significantly advances fine-grained understanding. By integrating advanced alignment techniques with large-scale, high-quality datasets and hard negative samples, FG-CLIP captures global-level and region-level semantic details and distinguishes subtle differences more effectively. Extensive experiments across diverse downstream tasks validate the model’s superior performance. In addition, we propose FgGRN as a unified dataset that combines high-quality region-specific annotations with challenging fine-grained negative samples, offering a valuable resource for advancing multimodal research. Looking ahead, exploring the integration of more sophisticated multimodal models and expanding dataset diversity will be crucial for pushing the boundaries of fine-grained understanding.

### v3

@@ Abstract
Contrastive Language-Image Pre-training (CLIP) excels in multimodal tasks such as image-text retrieval and zero-shot classification but struggles with fine-grained understanding due to its focus on coarse-grained short captions. To address this, we propose Fine-Grained CLIP (FG-CLIP), which enhances fine-grained understanding through three key innovations. First, we leverage large multimodal models to generate 1.6 billion long caption-image pairs for capturing global-level semantic details. Second, a high-quality dataset is constructed with 12 million images and 40 million region-specific bounding boxes aligned with detailed captions to ensure precise, context-rich representations. Third, 10 million hard fine-grained negative samples are incorporated to improve the model’s ability to distinguish subtle semantic differences. We construct a comprehensive dataset, termed FineHARD, by integrating high-quality region-specific annotations with hard fine-grained negative samples. Corresponding training methods are meticulously designed for these data. Extensive experiments demonstrate that FG-CLIP outperforms the original CLIP and other state-of-the-art methods across various downstream tasks, including fine-grained understanding, open-vocabulary object detection, image-text retrieval, and general multimodal benchmarks. These results highlight FG-CLIP’s effectiveness in capturing fine-grained image details and improving overall model performance. The data, code, and models are available at https://github.com/360CVGroup/FG-CLIP.
@@
Dataset: https://huggingface.co/datasets/qihoo360/FineHARD
@@ 1 Introduction
Third, to further enhance model robustness and discrimination abilities, we introduce a large-scale corpus of 10 million hard fine-grained negative samples. By incorporating these challenging negative samples into the training process, FG-CLIP learns to distinguish subtle differences in semantically similar but distinct pairs, thereby significantly improving its performance across various downstream tasks. We integrate the high-quality visual grounding data and hard fine-grained negative samples as a whole dataset called FineHARD.
@@ 3.2 Curated Dataset
Fine-Grained Visual Grounding+Recaption+Hard Negative Dataset (FineHARD).
For the second stage, we develop a high-quality visual grounding dataset named FineHARD, featuring precise region-specific captions and hard negative samples. We curate the overall dataset based on GRIT (Peng et al., 2024) images. The process begins with generating detailed image captions using CogVLM2-19B (Hong et al., 2024), ensuring comprehensive and nuanced descriptions that capture the full context of each image. Following (Peng et al., 2024), we then use SpaCy (Honnibal et al., 2020) to parse the captions and extract the referring expressions. Subsequently, the images and referring expressions are fed into the pretrained object detection model, i.e., Yolo-World (Cheng et al., 2024) to obtain the associated bounding boxes. Non-maximum suppression is applied to eliminate overlapping bounding boxes, retaining only those with predicted confidence scores higher than 0.4. This process results in 12 million images and 40 million bounding boxes with fine-grained region captions. We provide examples of the images and their corresponding captions in Appendix A.
@@ 4.1 Implementation Details
In the first stage, we train on a dataset of 1.6 billion images, each paired with short and long texts. The model is initialized with weights from the original CLIP (Radford et al., 2021). For both ViT-B and ViT-L (Dosovitskiy, 2021) configurations, the batch size per NPU is set to 384. The learnable temperature parameter $\tau$ is initialized to 0.07. We utilize the AdamW optimizer with a learning rate of 1e-4, weight decay of 0.05, $\beta_{1}$ of 0.9, $\beta_{2}$ of 0.98, and warmup steps for the first 200 iterations. The entire training process employs DeepSpeed’s Zero-2 optimization technique and Bfloat16 precision to accelerate training, and the model is trained for one epoch.
@@
In the second stage, we train on a dataset of 12 million images. Apart from long and short captions, this dataset includes high-quality visual grounding annotations and hard fine-grained negative samples. The model is initialized with weights obtained from the first stage. The batch size per GPU is set to 512. We employ the AdamW optimizer with a learning rate of 1e-6, weight decay of 0.001, $\beta_{1}$ of 0.9, $\beta_{2}$ of 0.98, and warmup steps for the first 50 iterations. Training acceleration techniques include DeepSpeed’s Zero-2 optimization, CUDA’s TF32 technology, and Bfloat16 precision, and the model is trained for one epoch.
@@ 5 Conclusion
In this work, we introduce Fine-Grained CLIP (FG-CLIP), a novel approach that significantly advances fine-grained understanding. By integrating advanced alignment techniques with large-scale, high-quality datasets and hard negative samples, FG-CLIP captures global-level and region-level semantic details and distinguishes subtle differences more effectively. Extensive experiments across diverse downstream tasks validate the model’s superior performance. In addition, we propose FineHARD as a unified dataset that combines high-quality region-specific annotations with challenging fine-grained negative samples, offering a valuable resource for advancing multimodal research. Looking ahead, exploring the integration of more sophisticated multimodal models and expanding dataset diversity will be crucial for pushing the boundaries of fine-grained understanding.